OpenStudy (anonymous):

Which expression is equivalent to 19x − 6x + 2 and why? 15x, because 19x − 6x + 2 = 13x + 2 = 15x 11x, because 19x − 6x + 2 = 19x − 8x = 11x 2 + 13x, because (19x − 6x) + 2 = 13x + 2 = 2 + 13x 2 − 13x, because (19x − 6x) + 2 = −13x + 2 = 2 − 13x

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

2 + 13x, because (19x − 6x) + 2 = 13x + 2 = 2 + 13x

OpenStudy (anonymous):

Which statement best describes the expression 3y ÷ 8? 8 divided by 3 times y 8 times y divided by 3 3 divided by 8 times y 3 times y divided by 8

OpenStudy (princeharryyy):

3 times y divided by 8

OpenStudy (anonymous):

Which of these is the algebraic expression for "4 less than the product of 5 and some number?" 5 − 4p 4 − 5p 5p − 4 4p − 5

OpenStudy (princeharryyy):

5p − 4

OpenStudy (anonymous):

Which statement best describes the expression 3 + y ÷ 2? The quotient of 2 and the sum of 3 and y The quotient of the sum of 3 and y, and 2 The sum of 3 and the quotient of 2 and y The sum of 3 and the quotient of y and 2

OpenStudy (princeharryyy):

The sum of 3 and the quotient of y and 2

OpenStudy (anonymous):

2 more? if you dont mind.

OpenStudy (princeharryyy):

ok

OpenStudy (princeharryyy):

quick

OpenStudy (anonymous):

ty so much harry your a life saver :D

OpenStudy (anonymous):

What is the value of fraction 1 over 2x3 + 5.2y when x = 2 and y = 3? 8.2 9.2 18.6 19.6

OpenStudy (princeharryyy):

what's 2x3

OpenStudy (anonymous):

6

OpenStudy (princeharryyy):

no! no! I guess! its the x^3

OpenStudy (anonymous):

then i have no idea. xD..

OpenStudy (princeharryyy):

is 2x3 + 5.2y like (2x3 + 5.2y)

OpenStudy (anonymous):

OH. then it would be 11.2y

OpenStudy (princeharryyy):

nope!

OpenStudy (anonymous):

wot..? then im clueless xD

OpenStudy (princeharryyy):

What is the value of fraction 1 over 2x3 + 5.2y when x = 2 and y = 3? 8.2 9.2 18.6 19.6

OpenStudy (princeharryyy):

wait! let me see

OpenStudy (anonymous):

9.2?

OpenStudy (princeharryyy):

the question isn't clear?

OpenStudy (anonymous):

oh ok i got the answer! ty harry!

OpenStudy (princeharryyy):

r u done?